# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

### v2

@@ A.2.2 Stationary Regressor
$=W_{-j,t}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)+\dfrac{\eta_{j,t-1}^{(1)}+\zeta_{j,t}-w_{j,t}}{\widehat{\varsigma}_{j}}$
$=W_{-j,t}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)+\dfrac{\eta_{j,t-1}^{(1)}+n^{-\tau}\phi_{j,t-1}}{\widehat{\varsigma}_{j}},$
where the first equality is by the definition of $\widehat{r}_{j,t}$ in (2.9), the second row applies the pseudo-true regression model (3.10) and the equality (A.45), the third row applies (A.55), and the fourth row applies (A.43). Then by the definition of $\mathcal{Z}_{0,j}$ in (A.84), we have the following decomposition
$\mathcal{Z}_{j}^{(1)}=\dfrac{n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}+\dfrac{n^{-1/2}\sum_{t=1}^{n}\left(\eta_{j,t-1}^{(1)}+n^{-\tau}\phi_{j,t-1}\right)u_{t}}{\widehat{\varsigma}_{j}\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}.$ (A.86)
We first bound the first term. By (A.60), we have $\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}1$ and thus
$\left|\dfrac{n^{-1/2}\sum_{t=1}^{n}u_{t}\widetilde{W}_{-j,t-1}^{\top}\left(\widetilde{\varphi}-\widetilde{\varphi}^{*}\right)}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\right|$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\|n^{-1/2}\sum_{t=1}^{n}D_{-j}^{-1}W_{-h,t-1}u_{t}\|_{\infty}\cdot\|D_{-j}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)\|_{1}$
@@
where the second step applies MS24’s (B.63), and Proposition 1 in this current paper.
@@
$\dfrac{n^{-1/2}\sum_{t=1}^{n}\eta_{j,t-1}^{(1)}u_{t}}{\text{$\sigma_{u}\widehat{\varsigma}_{j}\cdot\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}=\sqrt{\dfrac{{\rm var}(\eta_{j,t}^{(1)})}{\widehat{\varsigma}_{j}^{2}\cdot\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}\cdot\frac{n^{-1/2}\sum_{t=1}^{n}\eta_{j,t-1}u_{t}}{\sigma_{u}\sqrt{{\rm var}(\eta_{j,t}^{(1)})}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ (A.88)
Finally, note that $\phi_{j,t-1}$ is mildly integrated. Again by Kostakis et al., (2015, Lemma B4(ii)) we have $\sum_{t=1}^{n}\phi_{j,t-1}u_{t}=O_{p}\left(n^{(\tau+1)/2}\right),$ and thus
$n^{-1/2}\sum_{t=1}^{n}n^{-\tau}\phi_{j,t-1}u_{t}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0.$
@@
By (A.86), (A.87), (A.88), and (A.89), we deduce (A.84) when $j\in\mathcal{M}_{z}.$ This completes the proof of Theorem 1. ∎
@@
$\widehat{\omega}_{j}^{{\rm XD}}=\frac{\text{$\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}=\frac{\text{$\sqrt{\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}}$}}{\left|\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}\right|}$
where $\check{r}_{j,t}=\widehat{r}_{j,t}\cdot\widehat{\varsigma}_{j}.$
CASE I. When $j\in\mathcal{M}_{x}$, the weak convergence (A.70) implies $(n^{-(1+\tau)}\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1})^{-1}=O_{p}(1),$ and (A.76) and (A.28) implies $\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}=O_{p}(n^{1+\tau}).$ Then

### v3

@@ A.3.1 Technical Lemmas
$=W_{-j,t}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)+\dfrac{\eta_{j,t-1}^{(1)}-n^{-\tau}C_{\zeta}\phi_{j,t-1}}{\widehat{\varsigma}_{j}},$ (A.136)
where the first equality is by the definition of $\widehat{r}_{j,t}$ in (2.11), the second row applies the pseudo-true regression model (3.14) and the equality (A.84), and the third row applies (A.125). Then by the definition of $\mathcal{Z}_{0,j}$ in (A.131), we have the following decomposition
$\mathcal{Z}_{j}^{(1)}=\dfrac{n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}+\dfrac{n^{-1/2}\sum_{t=1}^{n}\left(\eta_{j,t-1}^{(1)}+n^{-\tau}\phi_{j,t-1}\right)u_{t}}{\widehat{\varsigma}_{j}\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}.$ (A.137)
We first bound the first term by
$\left|n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)\right|$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\|n^{-1/2}\sum_{t=1}^{n}D_{-j}^{-1}W_{-h,t-1}u_{t}\|_{\infty}\cdot\|D_{-j}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)\|_{1}$
@@
where the second step applies MS24’s (B.63), and Proposition 1 in this current paper. By (A.99), we have $\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}1$ and thus
$\left|\dfrac{n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\right|=o_{p}(1).$ (A.139)
@@
$\dfrac{n^{-1/2}\sum_{t=1}^{n}\eta_{j,t-1}^{(1)}u_{t}}{\text{$\sigma_{u}\widehat{\varsigma}_{j}\cdot\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}=\sqrt{\dfrac{{\rm var}(\eta_{j,t}^{(1)})}{\widehat{\varsigma}_{j}^{2}\cdot\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}\cdot\frac{n^{-1/2}\sum_{t=1}^{n}\eta_{j,t-1}^{(1)}u_{t}}{\sigma_{u}\sqrt{{\rm var}(\eta_{j,t}^{(1)})}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ (A.140)
Finally, note that $\phi_{j,t-1}$ is mildly integrated. Again by kostakis2015robust we have $\sum_{t=1}^{n}\phi_{j,t-1}u_{t}=O_{p}\left(n^{(\tau+1)/2}\right),$ and thus
$n^{-1/2}\sum_{t=1}^{n}n^{-\tau}\phi_{j,t-1}u_{t}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0.$ (A.141)
@@
By (A.137), (A.138), (A.140), and (A.142), we have
$\mathcal{Z}_{j}^{(1)}$ $=o_{p}(1)+\dfrac{n^{-1/2}\sum_{t=1}^{n}\eta_{j,t-1}^{(1)}u_{t}}{\text{$\sigma_{u}\widehat{\varsigma}_{j}\cdot\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}+o_{p}(1)$ (A.143)
$\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}0,$ (A.144)
which verifies (A.131) when $j\in\mathcal{M}_{z}.$ This completes the proof of Theorem 1. ∎
@@
$\widehat{\omega}_{j}^{{\rm XD}}=\frac{\text{$\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}=\frac{\text{$\sqrt{\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}}$}}{\left|\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}\right|}$ (A.145)
where $\check{r}_{j,t}=\widehat{r}_{j,t}\cdot\widehat{\varsigma}_{j}$ as defined in (A.106).
CASE I. When $j\in\mathcal{M}_{x}$, (A.118) and (A.64) implies
$\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}=\dfrac{\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}}{\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}\cdot\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm lvar}(e_{j,t})}{2C_{\zeta}},$ (A.146)
which implies $\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}=O_{p}(n^{1+\tau}).$ In addition, the weak convergence (A.112) implies that $\frac{1}{\left|\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}\right|}=O_{p}\left(\frac{1}{n^{1+\tau}}\right)$. Then